# Global Hash Tables Strike Back! An Analysis of Parallel GROUP BY Aggregation

### v1

@@ 4.2. Scaling
Figure 6. End-to-end evaluation of scaling behavior of fully concurrent aggregation methods vs. partitioned aggregation. Folklore* is used for the ticketing step for fully concurrent aggregation and both atomic and thread local updates are evaluated for the update step. Similar to Figure 4 and Figure 5, performance as throughput and the speedup factor relative to single-threaded execution are plotted against thread count.
We evaluate scaling from end to end (including both ticketing and partial aggregation update) in Figure 6. We test the atomic and thread local methods for fully concurrent aggregation and graph the results against those for partitioned aggregation using local preaggregation. The locked update approach is omitted because it performs strictly worse than atomics as a concurrent update method.
@@
In low cardinality datasets, we see that fully concurrent aggregation using thread local updates matches the performance of the partitioned approach, an impressive feat given that the partitioned workload does basically no extra work in the partition-wise aggregation stage since all keys fit in the local hash table during local aggregation stage.
@@
In the high cardinality case, we find that thread local aggregation has a clear advantage over partitioning across all data distributions. At this cardinality, the local hash table spills most of its entries, which causes the partitioning to have to aggregate each value twice (once locally and once partition-wise). For the non-heavy-hitter distribution, the atomic method also displays an advantage over partitioning. Surprisingly, thread local aggregation does not exhibit decreased throughput at high thread counts in the end-to-end benchmarks, despite doing so in the isolated update benchmarks from Section 3.2. This behavior is attributable to the fact that the actual update step is fast enough that even if scaling poorly, the ticketing step still dominates most of the runtime, as seen in the performance breakdowns in Section 4.3.
@@
In the pure insert workload, fully concurrent aggregation also exhibits superior performance at high thread count, but only when using atomic updates. Since there is no contention in this case on the underlying partial aggregations, using atomics creates very little overhead. However, the same issues with heavy hitters in the high cardinality case also hold true in the unique key case. Such a workload is not displayed in the graphs but is present in Table 2. Since thread local aggregation is now less feasible as an alternative to deal with skew, the combination of unique keys and high skew continues to present a significant performance challenge.
Table 2. Speedup relative to partitioned aggregation at different thread counts and workloads. Speedup is measured as, for a given data distribution and thread count, the latency of partitioned aggregation divided by the latency of the fully concurrent aggregation method. Speedups at parity with partitioned aggregation (0.9-1.1) are underlined and cases where there is clear superior performance by fully concurrent aggregation are bolded. In almost all cases, a form of fully concurrent aggregation achieves parity or better, with the only exception unique keys with heavy hitters at high thread counts. Thread local aggregation by itself achieves parity or better against partitioning for all workloads except at high thread count with unique keys.
Threads 1 4 16 32
Update Method Atomic Thread Local Atomic Thread Local Atomic Thread Local Atomic Thread Local
Low cardinality (uniform) 0.97 1.32 0.09 1.29 0.04 0.98 0.03 0.99
Low cardinality (zipfian) 0.93 1.28 0.07 1.25 0.03 1.01 0.02 1.07
Low cardinality (heavy hitter) 0.93 1.30 0.11 1.34 0.03 0.94 0.02 0.98
High cardinality (uniform) 1.77 1.83 1.32 1.37 1.54 1.64 1.79 1.77
High cardinality (zipfian) 1.90 2.01 1.31 1.41 1.55 1.70 1.97 1.82
High cardinality (heavy hitter) 1.38 1.49 0.99 1.21 0.53 1.44 0.34 1.31
Unique keys (uniform) 1.74 1.83 1.08 1.13 1.04 0.81 1.16 0.65
Unique keys (zipfian) 1.95 2.07 1.22 1.24 1.08 0.84 1.12 0.68
Unique keys (heavy hitter) 1.46 1.59 0.92 1.01 0.54 0.74 0.34 0.48
@@
Notably, the performance advantage of fully concurrent aggregation against partitioned aggregation does not come from its scaling behavior, with fully concurrent aggregation tending to have equal or marginally lower speedup factors at high thread counts as compared to partitioned aggregation. However, due to the significantly lower overhead of only inserting into a hash table a single time (rather than performing preaggregation and partition-wise aggregation), fully concurrent aggregation can achieve superior performance across many workloads. Table 2 shows the speedup of the fully concurrent aggregation techniques compared to the partitioned approach. Underlined values indicate performance parity (latency within a factor of 0.1) and bold values indicate clearly better performance compared with partitioning.
@@
As we discussed in Section 3.2, ideally the method for partial aggregate updates should be chosen based on workload. While this conclusion still holds true once considering end-to-end results, the picture improves for the thread local method, with the difference in performance narrowing between it and atomics in the high cardinality case. Of the results from Table 2, for its worst performing measured workload, the thread local method only runs about twice as long as partitioned. Similarly, the atomic method never runs more than twice as fast as the thread local method. Thus, the tail performance of thread local aggregation, though not ideal, is certainly not catastrophic and is potentially outweighed by the speedups obtained in the vast majority of workloads. Therefore, if implementers were to only choose one method for aggregation, they should choose fully concurrent aggregation with thread local updates. Such an implementation would achieve better performance or parity with our studied alternatives across the majority of workloads, and at its worst exhibit manageable amounts of performance degradation.
4.3. Latency Breakdown
@@
In order to explain the scaling behavior of fully concurrent aggregation, we time each stage of the aggregation process and plot the proportion of time spent in Figure 7. The initialization includes allocation of global data structures (the ticketing hash table and the global vector of atomic partial aggregates for atomic aggregation). For thread local aggregation, this stage does not include allocation of partial aggregate vectors because they are lazily initialized by each thread in the update stage. The materialization stage consists of the work to turn results into a columnar format that can be pushed to the next query operator, including the cost of merging thread local partial aggregates. The ticketing and update stages are as discussed in Section 3.
@@
Figure 7. Proportion of time spent on each step of fully concurrent aggregation. Ticketing consistently takes much more time than the update step except in high contention workloads atomic update workloads, which becomes a bottleneck. Materialization becomes a significant part of thread local aggregation’s runtime at high cardinality and thread count.
@@
We find that for fully concurrent aggregations, ticketing typically takes significantly more time than updating the partial aggregates. The exception to this trend is the low cardinality case for atomic updates, which at 4 threads becomes the major performance bottleneck and dominates runtime. Therefore, the choice of update step method should be driven more by tail performance. As long as the update step does not become a performance bottleneck, even relatively poor scaling in other steps is not insurmountable. This is the major factor that allows thread local updates to maintain relatively good performance even in its less ideal cases.
@@
Further, we find that the thread local method’s materialization cost becomes increasingly significant as thread count increases, which aligns with our experimental results and theoretical findings from Section 3.2: the work per thread of merging scales linearly with cardinality and is constant with regard to thread count. The hope was that the materialization step was fast enough in implementation that it would be manageable, which does hold true for even high cardinality workloads. However, once cardinality is so high that each key is unique, it becomes clear that the anti-scaling behavior of the update step is the significant factor contributing to degraded performance.
Memory Allocation
One surprising observation is that, for fully concurrent aggregations, a significant part of the time is spent on memory allocation, particularly as cardinality increases and thus the amount of memory needed for both the ticketing hash map and partial aggregate vectors. The surprisingly poor performance of the initialization stage suggests that performance gains to be had by optimizing the memory allocator for concurrent operations. We identify two major contributors to this issue: memory allocation for the ticketing step’s hash table and atomic updates’ partial aggregate vector is single-threaded, and construction of atomics is not always efficient (depending on language and compiler behavior, they may not be able to be zero-initialized). The former problem can probably be mitigated with techniques such as sharding, and the latter by leveraging more low-level language features. Database systems are also known to be sensitive to the performance and characteristics of the chosen memory allocator (Durner et al., 2019).
For the purposes of our experiments, we avoid over-tuning the memory allocation stage in order to more clearly isolate the impacts of the aggregation method itself, however based on the performance breakdowns this omission could have contributed to worse performance numbers for fully concurrent aggregation in the unique key workload, especially for atomic updates. However, since partitioned aggregation does not face the same memory allocation issues, these concerns do not conflict with our finding that fully concurrent aggregations are practical.
@@ 4.4. Resizing
A particularly thorny challenge with concurrent hash tables is resizing them efficiently. In most cases, resizing requires all other threads to pause work to accommodate reallocation and migration. While the hope is that cardinality estimates can allow a properly sized initial allocation that does not require resizing (Kipf et al., 2019), query optimizers are known to misestimate cardinality by many orders of magnitude under certain workloads (Leis et al., 2015).
Figure 8. Scaling behavior with a resize. Dashed lines indicate execution without resizing while solid lines indicate execution with resizing. Fully concurrent aggregation exhibits far more performance degradation under these circumstances while partitioned aggregation is not affected.
To test the impact of hash table resizing, we adopt Maier et al.’s (Maier et al., 2018) method for contention-less fully concurrent migration of hash table entries for Folklore*. In this experiment, we set the capacity of the ticketing hash table and partial aggregate vectors to be half of the required capacity, forcing a resize. We do similarly for the hash tables in partitioned aggregation (some ad-hoc configuration was required due to the standard library hash table’s allocator, but regardless, all workloads are tuned to result in exactly one resize). Figure 8 shows that the fully concurrent workload is particularly sensitive to resizing, which causes particularly high amounts of performance degradation at high thread count. While in the high cardinality workload, the fully concurrent methods still matche the performance of the partitioned workload, when the workload becomes insert-only (unique keys), resizing causes particularly severe performance degradation. Meanwhile, in the partitioned case, there is negligible impact from resizing. Thus, while not a show stopper, fully concurrent aggregation appears to be more sensitive to hash table resizing than partitioning approaches. Future work on improving the performance of such resizes has significant headroom for improvement.
4.5. Memory Usage
@@
Threads 1 8 32
Method Atomic Thread Local Partition Atomic Thread Local Partition Atomic Thread Local Partition
Low cardinality 0.001 0.001 0.001 0.004 0.004 0.008 0.016 0.018 0.031
High cardinality 0.522 0.522 4.501 0.526 1.047 3.913 0.537 2.849 3.480
Unique keys 5.216 5.216 9.038 5.219 10.435 9.038 5.231 28.330 9.088
Our model of execution assumes a main-memory system where all data structures fit within memory. It is not obvious how to adapt fully concurrent aggregation to disk spilling, unlike partitioning-based approaches like DuckDB’s (Kuiper et al., 2024). Therefore, understanding the severity of memory usage can be relevant for the feasibility of aggregation over large datasets.
In Section 4.5, we note that thread local updates create significant memory overhead while atomic updates are very memory-efficient. Analyzing the partitioned method is less clear, since much depends on the spilling behavior. However in the very worst case, when almost all keys are spilled, performance is bounded by the total number of elements (not just unique keys), which can severely degrade memory performance in even moderate cardinality workloads.
Comparing the peak memory usage of each technique in Table 3, we find that our theoretical findings hold. Atomic aggregation performs best, while partitioned aggregation displays very high memory overhead for higher cardinality cases. Thread local aggregation exhibits surprisingly good memory usage characteristics, using fewer memory resources than partitioned aggregation in all tested thread counts for the high cardinality workload, showing that the dense storage of partial aggregates can greatly counterbalance memory concerns. Even with unique keys, thread local aggregation’s memory usage is close to parity with partitioned aggregation at 8 threads. At 32 threads, however, memory usage degrades to be more 3.1x higher than that of partitioned aggregation and 5.4x that of the atomic update method. In any case, as we noted, at such high cardinalities it would be likely preferable to use atomic updates. Although memory considerations should be carefully considered on a system and workload basis, our results indicate that fully concurrent aggregation has significant advantages over partitioning across the majority of cases, although it is not obvious how to spill to disk when doing fully concurrent aggregations.

### v2

@@ 4.1. Scaling
[…]
End-to-end evaluation of scaling behavior of fully concurrent aggregation methods vs. partitioned aggregation. Folklore* is used for the ticketing step for fully concurrent aggregation and both atomic and thread local updates are evaluated. Similar to Figure 4 and Figure 6, performance as throughput and the speedup factor relative to single-threaded execution are plotted against thread count.
Figure 7. End-to-end evaluation of scaling behavior of fully concurrent aggregation methods vs. partitioned aggregation. Folklore* is used for the ticketing step for fully concurrent aggregation.
We evaluate scaling end-to-end (including both ticketing and partial aggregation update steps) in Figure 7. We test the atomic and thread local methods for fully concurrent aggregation and graph the results against those for partitioned aggregation using local preaggregation. The locked update approach is omitted because it performs strictly worse than atomics as a concurrent update method.
@@
At low cardinality, fully concurrent thread local aggregation outperforms the other methods at all thread counts, achieving 1.78x throughput compared to the partitioned aggregation at 48 threads. As in Section 3.2, atomic aggregation does not scale well.
@@
In the high cardinality case, we find that thread local aggregation has a clear advantage over partitioning at lower thread counts. At this cardinality, the local hash table used by partitioned aggregation spills most of its entries, which causes the partitioning to have to aggregate each value twice (once locally and once partition-wise). However, at high core counts the benefit narrows. For the non-heavy-hitter distribution, the atomic method also displays an advantage over partitioning, even in the presence of moderate skew in the Zipfian workload.
Thread local aggregation does not stop scaling until reaching above 32 threads, despite doing so in the isolated update benchmarks from Section 3.2. This behavior is because the update step is fast enough at low thread count such that the behavior of the ticketing step dominates the runtime, as seen in the performance breakdowns in Section 4.2.
Surprisingly, partitioned aggregation continues to scale when thread count exceeds the number of physical cores. However, even at the high range of thread counts, partitioned aggregation is around parity or slightly worse than fully concurrent methods. The exception is the heavy hitter case, where partitioned aggregation becomes a clear winner as a result of the performance gains from local preaggregation (which captures all the heavy hitter entries). However, this is a relatively extreme case of preaggregation effects.
@@
In the pure insert workload, fully concurrent aggregation also exhibits superior performance at high thread count, but only when using atomic updates. Thread local aggregation scales reasonably well at lower thread counts, but performance significantly degrades as the number of threads increases. Therefore, implementations using thread local updates should note to limit the maximum thread count under certain workload conditions. In this dataset, since there is no contention on the underlying partial aggregations, using atomics creates very little overhead.
@@
Notably, the performance advantage of fully concurrent aggregation against partitioned aggregation does not come from its scaling behavior, with fully concurrent aggregation achieving excellent scaling at high thread counts. However, fully concurrent aggregation has significantly lower hash table lookup overhead, since it only inserts into the ticketing hash table a single time (rather than performing aggregation locally and partition-wise).
In almost all cases, a form of fully concurrent aggregation achieves parity or better, with the only exception being high cardinality with heavy hitters at high thread counts due to the poor performance of atomic aggregation in the presence of skew. However, our heavy hitter case is a particularly extreme case of skew with 50% the same value. More realistic Zipfian-distributed workloads can be handled by atomic updates without too much issue.
Furthermore, the fact that thread local aggregation achieves its maximum speedup at a relatively low number of threads is advantageous in many real-world applications. It is important in practice to actually impose a max thread count for the thread local case, freeing up resources for other queries. Practitioners should consider if using twice as many resources (on partitioning) for a small lift in throughput is worthwhile, or if those additional resources are better allocated to more concurrent queries. We imagine that this tradeoff is acceptable for some, but not all, systems.
@@
As we discussed in Section 3.2, ideally the method for partial aggregate updates should be chosen based on workload. Given that in almost all cases, some form of fully concurrent aggregation achieves markedly superior performance over partitioning, a dynamic system could yield significant speedup over current practice. However, perhaps one of the more important findings of this paper is that no single aggregation method (including partitioning) is desirable in all cases. That being said, thread local aggregation shows significant versatility. It has acceptable tail behavior and superior throughput in many workloads. Furthermore, its excellent performance at low thread counts across all workloads makes it ideal for cases where resources are shared or limited. Therefore, if choosing only one aggregation method for a system, fully concurrent aggregation with thread local updates is a strong candidate. While we do not endorse one single aggregation method as the definitive choice for a database, this work on fully concurrent aggregation expands the toolkit available for designing a well-optimized query executor.
4.2. Explaining Behavior
In this section, we attempt to explain the scaling behavior of each aggregation method. We first take a high-level look by breaking down the time spent on each task (ticketing, partial aggregate update, etc). Then, we analyze specific performance counters for a deeper understanding of how execution characteristics change with thread count.
Task Breakdown
@@ 4.2. Explaining Behavior
First, we time each stage of the aggregation process and plot the proportion of time spent in Figure 8. The ticketing and update stages are as discussed in Section 3 and the preaggregation and partition-wise aggregation stages are as discussed in Section 2.2. The materialization stage consists of the work to turn results into a columnar format that can be pushed to the next query operator, including the cost of merging thread local partial aggregates and combining each partition’s aggregated values.
[…]
@@
Percent of time spent on each step of aggregation. Ticketing consistently takes much more time than the update step except in high contention workloads atomic update workloads, which becomes a bottleneck. Materialization becomes a significant part of thread local aggregation’s runtime at high cardinality and thread count.
Figure 8. Percent of time spent on each step of aggregation.
@@
We find that for fully concurrent aggregations, ticketing typically takes significantly more time than updating the partial aggregates. The exception is the low cardinality case for atomic updates, which at 4 threads becomes the major performance bottleneck. Therefore, the choice of update step method should be driven more by tail performance. As long as it does not become a performance bottleneck, even relatively poor scaling in the update step is not insurmountable. This is the primary factor that allows thread local updates to maintain relatively good performance.
@@
Further, we find that the thread local method’s materialization cost becomes increasingly significant as thread count increases, which aligns with our experimental results and theoretical findings from Section 3.2. Materialization is fast enough in many cases that its poor scaling is not a bottleneck. However, at high thread counts with very high cardinalities, the materialization step can become the driving factor for degraded performance.
The materialization effects of partitioned aggregation are also particularly interesting. The materialization of each partition-wise hash table and copying of values to the final vector of all outputs is parallelized. The decrease in materialization time at very high thread counts appears to contribute to the improved scaling of partitioned aggregation even after thread count exceeds real core count. We attribute some of this odd behavior to the size of the hash table of partition-wise aggregates, which must be walked in order to materialize the key and value output columns. The size of each partition’s hash table decreases with the number of partitions.
Performance Counters
[…]
Breakdown of time spent on work by processor per workload and thread count. Shows both idle time and topdown analysis of time performing work.
Figure 9. Breakdown of time spent on work by processor per workload and thread count. Shows both idle time and topdown analysis of time performing work.
[…]
Instruction count and IPC vs. threads. Contributes to understanding the performance characteristics of each workload and the amount of work to be performed.
Figure 10. Instruction count and IPC vs. threads.
We also gather performance counter data for each workload using Linux’s perf tool to observe how execution characteristics vary with thread count. Analysis of the resulting data is made difficult by the complicated nature of the workload, whose performance is driven by multiple factors, including contention, time complexity scaling, etc. We therefore take a layered approach and separate the analysis into three questions:
  1. (1)
    Q1: How much time is spent performing “work.” Specifically, we measure the proportion of time a worker thread is scheduled on a core. The compliment of this time is idle time, which occurs due to lock contention, synchronization effects, or lack of resources.
  2. (2)
    Q2: Of the time spent on “work,” what are the performance characteristics? These performance characteristics impact the rate that the processor performs work, that is instructions per cycle (IPC). We also employ topdown analysis (Yasin, 2014) for granular insights into the bottlenecks.
  3. (3)
    Q3: How much “work” is there? We measure the amount of work with total instruction count.
In Figure 9 we break down work time with topdown analysis as well as plot the split of time working and idling (Q1, Q2). In Figure 10 we plot the instructions and IPC (Q2, Q3).
At a low cardinality, we see that atomic updates quickly becomes backend bound (contention from atomic loads is considered a backend bound effect, not idle time, unlike lock contention). In contrast, the thread local and partitioned methods show relative stability in its execution characteristics until idle time and SMT contention start to become a factor. In addition, a high proportion of instructions are retiring, indicating that execution is compute-bound and not bottlenecked by memory, speculation effects, etc.
At higher cardinalities, more work is spent backend bound across the board. This behavior is expected given that at high cardinalities, every hash table lookup and partial aggregate update is expected to be a cache miss. We also see some idle time occurring at higher thread counts for the partitioned method, even prior to SMT. This observation points to a theoretical drawback of the approach. Partitioning tends to be more prone to synchronization issues because all threads must wait for the final partition to complete its task. Meanwhile, work-stealing is generally difficult in such a setup.
Now turning to the instruction and IPC counts in Figure 10, we confirm that the total work performed by thread local updates increases with thread count due to materialization. Also revealing is that while the base instruction count of partitioned aggregation stays stable, it is considerably higher than the other methods, owing to partitioning’s added overhead. The counterbalancing effect is its relatively higher IPC at higher cardinalities. This behavior helps to explain why SMT benefits partitioned aggregation in particular. The high instruction count and IPC indicates that the processor is better able to utilize the time spent stalling on memory or other resources to continue execution on other compute-bound work.
4.3. Other Platforms
In addition to the AMD platform, we also ran experiments on machines with ARM and Intel processors. The different hardware characteristics of each are described in Section 2.4. We graph the throughput achieved on these systems in Figure 11. We find that although the best technique for a given configuration may vary, the overall shape of the graph remains similar to our results in Figure 7.
On ARM, we find that partitioned aggregations performs particularly poorly on the high cardinality case. We found that on this system, a larger amount of time was spent on materialization, which combines all the thread local hash tables into two output columns of keys and values. This step is almost purely reads and writes to memory. Therefore, we believe this poor performance could relate to the lower memory bandwidth of our ARM setup, which is 3200MT/s as compared to 4800MT/s on our AMD setup and 5600MT/s on our Intel setup.
Note that the Intel chip has fewer cores, so the graphs do not reveal severe performance degradation for thread local aggregation at high thread counts. Here, the performance of thread local updates is excellent in all but the unique case, and even then performs admirably. Atomic updates, in turn, perform much better than partitioned at higher cardinalities. These results reinforce our discussion that on more resource-constrained machines, fully concurrent aggregation is an especially convincing choice.
[…]
Scaling behavior of aggregation on ARM (Ampere Altra Q80-30) and Intel (Xeon Gold 5412U) machines. Extended benchmarks and profiling results on skewed datasets can be found at rm.cab/ghplots.
Figure 11. Scaling behavior of aggregation on ARM (Ampere Altra Q80-30) and Intel (Xeon Gold 5412U) machines. Extended benchmarks and profiling results on skewed datasets can be found at rm.cab/ghplots.
4.4. Tuple Size
[…]
Throughput graphed against size in bits of value. For higher cardinality workloads, we see increasing the size of the aggregated value decreases throughput significantly for thread local aggregation, but marginally for the rest.
Figure 12. Throughput vs. size of value being aggregated.
Up until now, we assumed that all values were 64-bit integers. We now investigate the performance impact of the size of the tuple. Specifically, we hold the thread count (48) and key size (64-bit) constant and vary the value size to 16-bit, 32-bit, and 64-bit integers. We choose vary the output size because it impacts the update step, whereas the key size mostly influences the speed of hashing. We plot throughput against value size in Figure 12.
We find that the effect on performance is most pronounced for the higher cardinality workloads, which aligns with their backend bound behavior and thus sensitivity to factors such as memory bandwidth. The most significant impact is on thread local aggregation, which has clearly decreasing throughput at higher value sizes, especially in the unique keys case. This is because materialization of thread local aggregates requires many sequential reads from memory. The total number of reads during this operation increases with tuple size. On the other hand, atomic updates and partitioned aggregation have mostly random memory accesses, which incur a cache miss regardless of tuple size.
@@ 4.5. Resizing
A particularly thorny challenge with concurrent hash tables is how to resize them efficiently. In most cases, resizing requires all other threads to pause work to accommodate reallocation and migration. Although cardinality estimation can often allow a properly sized initial allocation (Kipf et al., 2019), poor estimates are always a possibility.
To test the impact of hash table resizing, we adopt Maier et al.’s (Maier et al., 2018) method for contention-less fully concurrent migration of hash table entries for Folklore*. In this experiment, we set the capacity of the ticketing hash table and partial aggregate vectors to be half of the required capacity, forcing a resize.
Figure 13 shows that the fully concurrent workload is does display significant performance degradation at higher thread counts in the presence of resizing. Resizing causes a considerable but not insurmountable latency increase for the high cardinality case. At 48 threads, atomics increase in latency by 1.8x and thread local by 1.6x. However, in the very worst case, unique keys, the degradation is much more extreme with a 5.2x latency increase for atomics and 2.0x for thread local. Thus, while not a show stopper, fully concurrent aggregation appears to be sensitive to the cost of resizing, especially in the most extreme cardinality cases. Future work on improving the performance of such resizes has significant headroom for improvement. Resizing performance and should be an important dimension for analyzing the suitability of concurrent hash table designs for database applications, and we hope this work motivates more research in the area.
[…]
Scaling behavior of fully concurrent aggregation with a resize. Dashed lines indicate throughput without resizing and solid lines with. Fully concurrent aggregation exhibits far more performance degradation under these circumstances. Partitioned approach is not dependent on accurate cardinality estimation, but it is plotted for comparison.
Figure 13. Scaling behavior of fully concurrent aggregation with resize. Dashed lines are without resizing and solid lines are with. Partitioned plotted for reference.
4.6. Memory
@@
Threads 1 8 48
Method Atomic Thread Local Partitioned Atomic Thread Local Partitioned Atomic Thread Local Partitioned
Low card. 0.001 0.001 0.000 0.004 0.004 0.004 0.024 0.028 0.024
High card. 0.299 0.299 2.898 0.302 0.824 3.133 0.322 3.828 2.971
Unique keys 4.098 4.098 4.625 4.102 9.317 5.000 4.121 39.143 5.137
Our model of execution assumes that all data structures fit within memory. It is not obvious how to adapt fully concurrent aggregation to disk spilling, unlike partitioning-based approaches (Kuiper et al., 2024). Therefore, understanding the severity of memory usage can be relevant for the feasibility of aggregation over large datasets.
In Section 3.2, we note that thread local updates create significant memory overhead while atomic updates are very space efficient. Analyzing the partitioned method is less clear, since much depends on the spilling behavior. However, in the very worst case, when almost all keys are spilled, memory usage is bounded by the total number of elements (not just unique keys), which can severely increase memory usage.
Comparing the peak memory usage of each technique in Table 3, we find that our theoretical model holds. Atomic aggregation performs best, while partitioned aggregation displays very high memory overhead for higher cardinality cases. Thread local aggregation exhibits surprisingly good memory usage characteristics, using much fewer memory resources than partitioned aggregation until the high end of thread count. In the unique case, at very high thread counts, thread local aggregation has much higher memory usage. However, since we would limit the number of threads in practice, the memory usage would not reach the extreme case. Although memory considerations should be carefully considered on a system and workload basis, our results indicate that fully concurrent aggregation has substantial advantages over partitioning across many cases, although it is not obvious how to spill to disk.